OpenStudy (anonymous):

HELP PLEASE!! Jennifer wants to take a trip around the world. She plans to deposit $125 at the beginning of each month into an investment with a 3.75% interest rate, compounded monthly. How much will she have in the account after 8 years? $13,969.10 $14,012.75 $110,885.27 $115,043.47

OpenStudy (beccaboo333):

First you want to find how much the interest is that will be added.

OpenStudy (anonymous):

how do i do that? @beccaboo333

OpenStudy (beccaboo333):

I confused myself @jdoe0001 @dumbcow @iambatman

OpenStudy (dumbcow):

this involves using a financial calculator or a somewhat complicated formula the formula is based on the sum of a geometric series \[S = 125 \frac{r (r^{96} -1)}{r-1}\] where \[r = 1 + \frac{i}{12}\]

OpenStudy (anonymous):

it has to do with future value i just dont understand how to do it @beccaboo333

OpenStudy (dumbcow):

i know that is the future value equation

OpenStudy (anonymous):

so what is FV= PV( 1+i)nt i thought that was future value formula

OpenStudy (dumbcow):

ahh it is but we don't know the present value either all we know is she is depositing 125 every month for 96 months

OpenStudy (anonymous):

im kind of confused i thought 125 would be the present value because thats whats being deposited

OpenStudy (dumbcow):

it would be if she just left it there, but every month she is adding another 125, this makes it more complicated

OpenStudy (anonymous):

so in the formula you wrote does r represent the interest rate?

OpenStudy (dumbcow):

r is the growth factor i is annual int rate